OpenStudy (anonymous):

Solve the equation. Be sure to show all of your work. 10n – 9 = 61

OpenStudy (anonymous):

add 9 divide by 10

OpenStudy (anonymous):

\[10n-9=61\]add 9 \[10n-9+9=61+9\] \[10n=70\] divide by 10\[\frac{10n}{10}=\frac{70}{10}\] \[n=7\]